# Identifying treatment response subgroups in observational time-to-event data

### v4

@@ B.5.3 Impact of treatment rate
B.5.4 Aligned covariates and treatment effect structure
Section B.1 presents a data generation with the last 8 covariates having a clustered structures independent of the treatment response. This section explores a setting where all clusters in the covariates are associated with different treatment responses of interest. Specifically, we sample the last dimensions from standard normal distributions instead of various clusters, i.e.,
@@ B.5.4 Aligned covariates and treatment effect structure
Even in this unrealistic scenario, our proposed method remains the second best in recovering the underlying clustering structure and associated treatment responses. This observation reinforces the main findings of our work, demonstrating the method’s potential to identify subgroups of treatment effects across different and even adverserial settings.
Appendix C Case study: Heterogeneity of adjuvant radiotherapy responses in the Seer dataset
To study the medical relevance of the proposed method, we explore data from the Surveillance, Epidemiology, and End Results program [^8] Available at https://seer.cancer.gov/ (Seer) gathering patients diagnosed with breast cancer between 1992 and 2017. Following (Lee et al., 2018; Danks & Yau, 2022; Jeanselme et al., 2023), we select women who died from the condition or from cardiovascular diseases. From this observational data, we investigate the impact of adjuvant radiotherapy after chemotherapy on survival outcomes. To this end, we subselect patients with recorded treatment and who received chemotherapy. These criteria led to the selection of 239,855 patients with 22 covariates measured at diagnosis such as diagnosis year, grades, ethnicity, laterality, tumour size and type (see  (Danks & Yau, 2022) for further description).
Selection of $K$.
From the selected population, our aim is to identify heterogeneous responses to adjuvant treatment. The first challenge is the selection of the number of groups to use ($K$). We advise to follow medical actionability and consider the change in treatment effects and size of the subgroups when increasing this parameter. In the absence of experts’ intuition, one may rely on an elbow rule heuristic over $l_{F}$. Using Figure 6, the negative log-likelihood presents an elbow for $K=2$.
@@ Appendix C Case study: Heterogeneity of adjuvant radiotherapy responses in the Seer dataset
Treatment effect subgroups.
In this section, we examine the treatment response following adjuvant radiotherapy to identify groups of patients who received surgery and chemotherapy and may benefit from adjuvant radiotherapy. This problem is central to patients’ treatment as no evidence-based guidelines for adjuvant therapy exist (Lazzari et al., 2023), making this setting more likely to meet the positivity assumption (Assumption 4), necessary to study causality in observational data.
(a) Causal Survival Clustering
(b) Cox Mixture of Heterogeneous Effects
@@
Figure 8: Causal Survival Clustering - Change in log-likelihood given random permutation of a given covariates.
Uncovering treatment response.
Figure 7 presents the identified treatment effect subgroups when using CSC and CMHE. For this latter, the number of subgroups $K$ is selected through hyperparameter tuning, leading to $K=1$ in every fold. As shown by Figure 6, using the previously described elbow rule leads to the same number of subgroups. As previously mentioned, our proposed methodology presents two strengths that explain the difference in the identified subgroups of treatment effects. First, the survival distribution under treatment is not constrained by the one under the control regime, resulting in more flexible, non-proportional distributions. CMHE’s parametrisation, which characterises treatment as a linear shift in the log hazard, results in a proportionality assumption between treated and untreated distributions. Second, CMHE does not account for treatment non-randomisation in its average treatment effect estimation, whereas our use of inverse propensity weighting corrects for any observed ones.
Using a permutation test, we identify as the covariates that most impact the likelihood associated with the model. Figure 8 displays the 10 covariates most indicative of the different treatment response subgroups.
Table 7 summarises the average value across the identified subgroups and the life expectancy gain when using adjuvant radiations measured through the Restricted Mean Survival Time (RMST) (Royston & Parmar, 2013). Both methodologies identify a population with limited treatment response. However, our proposed methodology identifies a second group, characterised by larger HER2 and larger distant lymph node count, with a positive treatment response, gaining more than half a year of life expectancy over the five years following diagnosis.
The proposed analysis pinpoints a group that could benefit from adjuvant radiotherapy. However, our methodologies remain hypothesis-generating tools, requiring further experimental validation, particularly due to potential confounding through hormonal therapy (not available in this dataset), the temporal nature of treatment, and the plurality of treatment options. The quality of available covariates limits this analysis and only serves as an example to medical practitioners to identify subgroups of treatment responses from observational data.

### v5

@@ C.5.3 Impact of treatment rate
Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ 25% CSC 0.830 (0.047) 0.051 (0.008) 0.023 (0.013) 0.020 (0.005) CSC Unadjusted 0.857 (0.033) 0.045 (0.006) 0.030 (0.015) 0.015 (0.003) CMHE ($L=3$) 0.376 (0.028) 0.168 (0.007) 0.077 (0.002) 0.073 (0.004) CMHE ($M=3$) 0.134 (0.090) 0.208 (0.028) 0.090 (0.015) 0.134 (0.030) CMHE ($M=L$) 0.410 (0.038) 0.223 (0.011) - - KMeans + TE 0.001 (0.006) 0.187 (0.005) 0.088 (0.014) 0.149 (0.011) Virtual Twins 0.527 (0.060) 0.034 (0.012) 0.019 (0.005) 0.034 (0.012) 75% CSC 0.718 (0.198) 0.061 (0.050) 0.057 (0.048) 0.023 (0.005) CSC Unadjusted 0.804 (0.045) 0.037 (0.006) 0.023 (0.008) 0.025 (0.007) CMHE ($L=3$) 0.385 (0.039) 0.168 (0.007) 0.077 (0.002) 0.073 (0.004) CMHE ($M=3$) 0.318 (0.140) 0.208 (0.028) 0.090 (0.015) 0.134 (0.030) CMHE ($M=L$) 0.410 (0.038) 0.223 (0.011) - - KMeans + TE 0.001 (0.006) 0.187 (0.005) 0.088 (0.014) 0.149 (0.011) Virtual Twins 0.527 (0.060) 0.034 (0.012) 0.019 (0.005) 0.034 (0.012)
Table 6: Cross-validated performance (with standard deviation in parentheses) with varying treatment rates under observational settings.
C.5.4 Aligned covariates and treatment effect structure
Sec. C.1 presents a data generation with the last 8 covariates having a clustered structure independent of the treatment response. This section explores a setting where all clusters in the covariates are associated with different treatment responses of interest. Specifically, we sample the last dimensions from standard normal distributions instead of various clusters, i.e.,
@@ C.5.4 Aligned covariates and treatment effect structure
Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ Randomised CSC 0.559 (0.114) 0.024 (0.009) 0.030 (0.017) 0.030 (0.006) CSC Unadjusted 0.481 (0.069) 0.026 (0.005) 0.041 (0.026) 0.032 (0.009) CMHE ($L=3$) 0.425 (0.020) 0.059 (0.004) 0.034 (0.025) 0.134 (0.012) CMHE ($M=3$) 0.165 (0.107) 0.062 (0.004) 0.126 (0.175) 0.233 (0.179) CMHE ($M=L$) 0.254 (0.138) 0.074 (0.027) 0.052 (0.021) 0.149 (0.018) KMeans + TE 0.888 (0.020) 0.015 (0.007) 0.021 (0.007) 0.020 (0.003) Virtual Twins 0.200 (0.090) 0.035 (0.011) 0.065 (0.025) 0.057 (0.030) Observational CSC 0.584 (0.151) 0.015 (0.007) 0.028 (0.024) 0.035 (0.003) CSC Unadjusted 0.556 (0.249) 0.023 (0.008) 0.051 (0.025) 0.040 (0.006) CMHE ($L=3$) 0.392 (0.062) 0.059 (0.008) 0.034 (0.019) 0.131 (0.015) CMHE ($M=3$) 0.133 (0.027) 0.059 (0.003) 0.051 (0.007) 0.153 (0.004) CMHE ($M=L$) 0.293 (0.125) 0.063 (0.012) 0.051 (0.017) 0.140 (0.011) KMeans + TE 0.895 (0.021) 0.013 (0.005) 0.042 (0.010) 0.020 (0.004) Virtual Twins 0.226 (0.069) 0.045 (0.024) 0.042 (0.031) 0.048 (0.011)
Table 7: Cross-validated performance (with standard deviation in parentheses) when all covariate clusters present different treatment responses. This setting aligns with the assumptions made by the step-wise approaches.
@@
Even in this unrealistic scenario, our proposed method remains the second best in recovering the underlying clustering structure and associated treatment responses. This observation reinforces the main findings of our work, demonstrating the method’s potential to identify subgroups of treatment effects across different and even adversarial settings.
Appendix D Additional results: Heterogeneity of adjuvant radiotherapy responses in the Seer dataset
This section presents additional results on the heterogeneity of adjuvant radiotherapy responses in the Seer dataset.
D.1 Selection of $K$
Using Figure 7, one can use the elbow rule on the negative log-likelihood. Following this heuristic, our analysis uses $K=2$ subgroups.
@@ D.1 Selection of $K$
D.2 Factors impacting subgroup membership
Figure 8: Causal Survival Clustering - Change in log-likelihood given random permutation of a given covariate.
Using a permutation test, we identify the covariates that most impact the likelihood associated with the proposed CSC model. Figure 8 displays the 10 covariates which most impact the likelihood, and therefore are associated with the different treatment response subgroups. This proposed analysis provides a tool to validate the model’s medical relevance by studying which covariates are associated with outcomes.
D.3 Baselines’ identified subgroups
[…]
@@ D.3 Baselines’ identified subgroups
Our main analysis describes how to use the proposed methodology to study a medical application. In this section, we analyse the groups one would identify with the previously described baselines. Figure 9 displays the identified treatment effect subgroups using all considered baselines. For CMHE, the number of subgroups $K$ is selected through hyperparameter tuning, leading to $K=1$ in every fold. As previously mentioned, our proposed methodology presents two strengths that explain the difference in the identified subgroups of treatment effects compared to CMHE. First, the survival distribution under treatment is not constrained by the one under the control regime, resulting in more flexible, non-proportional distributions. CMHE’s parametrisation, which characterises treatment as a linear shift in the log hazard, results in a proportionality assumption between treated and untreated distributions. Second, CMHE does not account for treatment non-randomisation in its average treatment effect estimation, whereas our use of inverse propensity weighting corrects for any observed ones. While both methodologies identify a group with limited response, our proposed methodology distinguishes a second group with improved treatment response.
Additionally, the KMeans approach identifies clusters that resemble those found with CSC, though the detected positive response is comparatively smaller. In contrast, the virtual twins method identifies a subgroup that exhibits a larger treatment effect, while the majority shows a negative response to treatment. Importantly, both methods broadly corroborate our findings. However, CSC presents a key strength by jointly learning group membership and treatment effects, ensuring that the covariates defining each subgroup are directly associated with the outcome – a property that two-step approaches cannot guarantee.